In a survey of 1000 US adults, 700 think police officers should be required to wear body cameras while on duty. Construct a 99% confidence interval for the population proportion.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 700 / 1000 = 0.700

1 - \hat p = 1 - 0.700 = 0.3

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.700 * 0.3) / 1000)

Margin of error = E = 0.037

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.700 - 0.037 < p < 0.700 + 0.037

0.663 < p < 0.737

The 99% confidence interval for the population proportion p is : 0.663 , 0.737
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 700 / 1000 = 0.700

1 - \hat p = 1 - 0.700 = 0.3

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.700 * 0.3) / 1000)

Margin of error = E = 0.037

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.700 - 0.037 < p < 0.700 + 0.037

0.663 < p < 0.737

The 99% confidence interval for the population proportion p is : 0.663 , 0.737